OpenStudy (anonymous):

Landon opened a savings account with a deposit of $3,786.45. Nine years later, the balance in the account is $5,135.29. If the interest rate is 3.4%, how often does the interest compound?

OpenStudy (anonymous):

why did it go away?

OpenStudy (akashdeepdeb):

I deleted it! I was on the wrong track! :D O.O

OpenStudy (anonymous):

Haha okayy.

OpenStudy (akashdeepdeb):

Tag someone!

OpenStudy (anonymous):

how?

OpenStudy (akashdeepdeb):

@ 'And then their username'

OpenStudy (amistre64):

a good formula to use would be the compounding formula: A = P(1+r/n)^(nt)

OpenStudy (amistre64):

A = Pk^(9n) A/P = k^(9n) ln(A/P) = 9n ln(k) ln(A/P)/9ln(k) = n ln(A/P)/9ln(1+r/n) = n this at least gives us a resurrsion setup that we can trial and error by making a guess n = ln(A/P)/9ln(1+r/g) ; each new value for n the we get we plug in got "g" and recalculate ... it should get us closer and closer to the true value of n

OpenStudy (amistre64):

of $3,786.45. Nine years later, the balance in the account is 5135.29. If the interest rate is 3.4%, h ln(5135.29/3786.45)/(9ln(1+.034/g)) i guessed at 12 and it crept down i guessed at 2 and it crept up i guessed at 6 and it crept down it has to be between 2 and 6 i guessed at 4 and it barely creeping .... lets assume 4 is good

OpenStudy (amistre64):

3786.45(1+.034/4)^(9*4) = 5135.2869...

OpenStudy (amistre64):

i spose we could have trial and errored it without the formula ... but you know :)